OpenStudy (anonymous):

How do I do this? 3square root tv multiplied by square root t

OpenStudy (anonymous):

\[\sqrt[3]{tv}\]

OpenStudy (anonymous):

\[\sqrt[3]{tv}\times \sqrt{tv}\]